Find the equation of the circle that passes through A(-3,1) and is concentric the circle x^2+y^2-3x+4y-10=0
Answer by richwmiller(17219)

Hint : Concentric circles will have the same center but a different radius.
The equation can be written (x-3/2)^2+(y+2)^2=65/4
radius is sqrt(65/4)
center is (1.5,-2)
(x-3/2)^2+(y+2)^2=
plug in (-3,1)
and find the new radius